OpenStudy (anonymous):

GIVING MEDALS!!! Isosceles Trapezoid ABCD has diagonals AC=10x+7 and BD = 2x+41. Find The Value Of x. :D

jigglypuff314 (jigglypuff314):

the diagonals in an Isosceles Trapezoid would be equal therefore AC = BD and so solve for x when 10x + 7 = 2x + 41

OpenStudy (anonymous):

4.5?

jigglypuff314 (jigglypuff314):

close, I'm getting 4.25

OpenStudy (anonymous):

No wait.

OpenStudy (anonymous):

Yeah, I got it :D

OpenStudy (anonymous):

If one base angle of an isosceles triangle measures 42 Degrees, what is the measure of the other base angle? :D

OpenStudy (anonymous):

42? Correct?

jigglypuff314 (jigglypuff314):

Correct! :D

jigglypuff314 (jigglypuff314):

the parallel line transversal thing (I forgot the name of it) 

OpenStudy (anonymous):

?

jigglypuff314 (jigglypuff314):

because of that rule, you can get that (5x + 13) + (3x - 5) = 180

OpenStudy (anonymous):

ik xD I got it :P

OpenStudy (anonymous):

21.5?

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

Thank you? Can You Do more?

jigglypuff314 (jigglypuff314):

of course :)

OpenStudy (anonymous):

If the length of one base of a trapezoid is 5, the length of the other base is 2x + 7, and the length of the midsegment is 6x-24, what is the value of x?

jigglypuff314 (jigglypuff314):

the midsegment is the average of the two bases so \[\frac{ (5) + (2x + 7) }{ 2 } = (6x - 24)\]

OpenStudy (anonymous):

I don't know how to find that :/

jigglypuff314 (jigglypuff314):

multiply both sides by 2 so (5 + 2x + 7) = 2(6x - 24) then distribute

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

If the length of one base of a trapezoid is 5, the length of the other base is 2x + 7, and the length of the midsegment is 6x-24, what is the value of x?

jigglypuff314 (jigglypuff314):

I've already gone over the first few parts of it above... solve for x when 2x + 12 = 12x - 48

OpenStudy (anonymous):

x=6? @jigglypuff314

jigglypuff314 (jigglypuff314):

sorry it took me so long, I had a class to go to, x = 6 would be correct :)